# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v1

@@ 3.2 Gradient Balancing Methods
Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ denoting the gradient of the $i$-th objective at the $k$-th iteration. This type of approach aims to find a common update direction ${\bm{d}}^{(k)}$ by adaptively aggregating the gradients of all objectives $\{{\bm{g}}_{i}^{(k)}\}_{i=1}^{m}$ at each iteration. Then, the model parameter is updated via ${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta{\bm{d}}^{(k)}$, where $\eta$ is a step size. To simplify notations, we omit the superscript in this section. Compared with loss balancing methods in Section 3.1, gradient balancing approaches usually achieve better performance.
@@ 3.2.1 Gradient Weighting Methods
where ${\bm{G}}=[{\bm{g}}_{1},\cdots,{\bm{g}}_{m}]\in\mathbb{R}^{d\times m}$ is the gradient matrix and ${\bm{\lambda}}=(\lambda_{1},\cdots,\lambda_{m})^{\top}$ is the gradient weights.
Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [64], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses, formulated as follows:
$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))$
$\approx$ $\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
where $\eta$ is the step size. The solution of problem (12) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where ${\bm{\lambda}}$ is computed as follows:
@@
Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a diagram $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant. Specifically, ${\bm{d}}$ is computed by solving the following problem:
@@
Problem (14) is equivalent to first compute ${\bm{\lambda}}$ by solving the following problem:
@@
where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate ${\bm{d}}$ as follows:
${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}.$ (16)
Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA [6] when $c\rightarrow+\infty$.
Similar to CAGrad [19], Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) [20] also constraints the update direction ${\bm{d}}$ stays not far away from a target direction (such as ${\bm{g}}_{0}$), while SDMGrad achieves it by leveraging an inner-product regularization ${\bm{g}}_{0}^{\top}{\bm{d}}$.
@@
${\bm{\lambda}}_{(2,\cdots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (18)
where ${\bm{\lambda}}_{(2,\cdots,m)}=(\lambda_{2},\cdots,\lambda_{m})^{\top}$, ${\bm{U}}=[{\bm{u}}_{1}-{\bm{u}}_{2},\cdots,{\bm{u}}_{1}-{\bm{u}}_{m}]$, and ${\bm{D}}=[{\bm{g}}_{1}-{\bm{g}}_{2},\cdots,{\bm{g}}_{1}-{\bm{g}}_{m}]$.
Nash-MTL [21] formulates problem (11) as a bargaining game [73, 74], where each objective is cast as a player and the utility function for each player is defined as ${({\bm{g}}_{i})}^{\top}{\bm{d}}$. Hence, ${\bm{d}}$ is computed by solving the following problem:
@@
FairGrad [22] then extends the more general formulation of Nash-MTL [21]. It computes ${\bm{\lambda}}$ by solving the following problem:
${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma},$ (21)
where $\gamma\geq 0$ is a constant. Different from Nash-MTL [21], FairGrad considers problem (21) as a simple constrained nonlinear least square problem.
Aligned-MTL [23] considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of ${\bm{G}}$. Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed. Then, the eigenvalues $\{\sigma_{1},\cdots,\sigma_{R}\}$ and eigenvectors ${\bm{V}}$ of ${\bm{C}}$ are obtained. Finally, ${\bm{\lambda}}$ is computed by:
@@
where ${{\bm{\Sigma}}}^{-1}=\text{diag}\left(\sqrt{\nicefrac{{1}}{{\sigma_{1}}}},\cdots,\sqrt{\nicefrac{{1}}{{\sigma_{R}}}}\right)$.
Gradient Normalization (GradNorm) [24] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have a similar magnitude. It updates ${\bm{\lambda}}$ by solving the following problem via one-step gradient descent:
@@
where $c=\frac{1}{m}\sum_{i=1}^{m}\left\|\lambda_{i}{\bm{g}}_{i}\right\|$ is the average scaled gradient norm and is cast as a constant, $r_{i}=\frac{f_{i}/f^{(0)}_{i}}{\frac{1}{m}\sum_{j=1}^{m}(f_{j}/f_{j}^{(0)})}$ denotes the training speed of the $i$-th objective and is also considered as a constant, and $\gamma>0$ is a hyperparameter.
@@ 3.2.2 Gradient Manipulation Methods
Gradient manipulation methods usually correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ to address the conflicting gradient problem and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
Projecting Conflicting Gradients (PCGrad) [25] defines that two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ are conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$. Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients. Specifically, at each iteration, $\hat{{\bm{g}}}_{i}$ is initialized with its original gradient ${\bm{g}}_{i}$. Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as follows:
@@
This process reduces gradient conflicts by removing the components of $\hat{{\bm{g}}}_{i}$ that oppose other objective gradients. Finally, the update direction is calculated as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
@@
where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$. $\bar{\phi}_{ij}$ is initialized with 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter. Note that GradVac is simplified to PCGrad if $\bar{\phi}_{ij}=0$.
@@ 3.2.3 Practical Speedup Strategy
Gradient balancing methods suffer from high computational and storage costs. Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$, which can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters. Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], SDMGrad [20], Nash-MTL [73], and FairGrad [22]) need to solve an optimization problem to obtain objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs. Hence, several strategies are proposed to alleviate this problem.
Sener and Koltun [6] use feature-level gradients (i.e., the gradient w.r.t the representations from the last shared layer) to replace parameter-level gradients (i.e., the gradient w.r.t shared parameters ${\bm{\theta}}$). Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs. This strategy is adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [23].
Liu et al. [19] randomly select a subset of objectives to calculate the update direction in each iteration, which is adopted by SDMGrad [20]. Navon et al. [21] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iteration instead of in each iteration. Although this strategy speedups the training process, they observe that it may cause a noticeable drop in performance. Liu et al. [75] consider optimizing the logarithm objective of MGDA [6] and propose a speedup strategy for it. Specifically, when solving problem (13) via one-step gradient descent, ${\bm{\lambda}}$ is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$. Note that,
$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$ (26)
$\approx\frac{1}{\eta}\left(f^{(k)}_{1}-f^{(k+1)}_{1},\cdots,f^{(k)}_{m}-f^{(k+1)}_{m}\right)^{\top},$ (27)
@@ 3.3 Discussions
Loss balancing and gradient balancing methods can effectively mitigate conflicts among different objectives and achieve a Pareto optimal solution, which is important for MOO scenarios such as multi-task learning. However, they are limited to finding a single Pareto optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning. In Sections 4 and 5, we introduce some methods to achieve more diverse and controllable solutions.
@@ 4 Finding a finite set of solutions
In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs. This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the PS. We first discuss preference vector-based methods in Section 4.1, followed by preference vector-free approaches in Section 4.2.
4.1 Preference Vector-Based Methods
Generally, preference vector-based methods rely on a preference vector set $\{{\bm{\alpha}}^{(1)},\cdots,{\bm{\alpha}}^{(n)}\}$. These preference vectors partition the problem into $n$ subproblems, where each subproblem involves finding the solution that corresponds to a specific preference vector in the set. By solving $n$ subproblems, these methods obtain a set of $n$ solutions.
Pareto Multi-Task Learning (PMTL) [28] incorporates preference vectors as constraints on the objectives. For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to other preference vectors, formulated as:
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=(f_{1}({\bm{\theta}}),\cdots,f_{m}({\bm{\theta}}))^{\top},$

### v2

@@ 3.2 Gradient Balancing Methods
Let ${\bm{g}}_{i}^{(k)}=\nabla_{\bm{\theta}}f_{i}({\bm{\theta}})|_{{\bm{\theta}}^{(k)}}\in\mathbb{R}^{d}$ be the gradient of the $i$-th objective at the $k$-th iteration. This type of approach aims to find a common update direction ${\bm{d}}^{(k)}$ by adaptively aggregating the gradients of all objectives $\{{\bm{g}}_{i}^{(k)}\}_{i=1}^{m}$ at each iteration. Then, the model parameter is updated via ${\bm{\theta}}^{(k+1)}={\bm{\theta}}^{(k)}-\eta{\bm{d}}^{(k)}$, where $\eta$ is a step size. To simplify notations, we omit the superscript in this section. Compared with loss balancing methods in Section 3.1, the gradient balancing approach usually achieves better performance.
@@ 3.2.1 Gradient Weighting Methods
where ${\bm{G}}=[{\bm{g}}_{1},\dots,{\bm{g}}_{m}]\in\mathbb{R}^{d\times m}$ is the gradient matrix and ${\bm{\lambda}}=[\lambda_{1},\dots,\lambda_{m}]^{\top}$ is the gradient weights.
Sener and Koltun [6] apply Multiple-Gradient Descent Algorithm (MGDA) [73], which aims to find a direction ${\bm{d}}$ at each iteration to maximize the minimal decrease across the losses. This is formulated as:
$\max_{{\bm{d}}}\min_{i\in[m]}(f_{i}({\bm{\theta}})-f_{i}({\bm{\theta}}-\eta{\bm{d}}))\approx\max_{{\bm{d}}}\min_{i\in[m]}{\bm{g}}^{\top}_{i}{\bm{d}},$ (12)
where $\eta$ is the step size. The solution of problem (12) is ${\bm{d}}={\bm{G}}{\bm{\lambda}}$, where ${\bm{\lambda}}$ is computed as:
@@
Conflict-Averse Gradient descent (CAGrad) [19] improves MGDA [6] by constraining the aggregated gradient ${\bm{d}}$ to be around the average gradient ${\bm{g}}_{0}=\frac{1}{m}\sum_{i=1}^{m}{\bm{g}}_{i}$ with a distance $c\left\|{\bm{g}}_{0}\right\|$, where $c\in[0,1)$ is a constant. Specifically, ${\bm{d}}$ is computed by solving the following problem:
@@
Problem (14) is equivalent to first computing ${\bm{\lambda}}$ by solving the following problem:
@@
where ${\bm{g}}_{\bm{\lambda}}=\frac{1}{m}{\bm{G}}{\bm{\lambda}}$ and then calculate the update direction ${\bm{d}}={\bm{g}}_{0}+\frac{c}{\left\|{\bm{g}}_{\bm{\lambda}}\right\|}{\bm{g}}_{\bm{\lambda}}$. Note that CAGrad is simplified to EW when $c=0$ and degenerates into MGDA [6] when $c\rightarrow+\infty$.
MGDA [6] and its variant CAGrad [19] can converge to a Pareto stationary point in the deterministic setting with full-gradient computations. However, in deep learning, where stochastic gradients are used, many works [20, 21, 23, 22, 24, 25, 26] explore the convergence properties of MGDA under stochastic gradients, which are introduced in detail in Section 6.1.
@@
${\bm{\lambda}}_{(2,\dots,m)}={\bm{g}}_{1}^{\top}{\bm{U}}\left({\bm{D}}{{\bm{U}}}^{\top}\right)^{-1},\quad\lambda_{1}=1-\sum_{i=2}^{m}\lambda_{i},$ (17)
where ${\bm{\lambda}}_{(2,\dots,m)}=[\lambda_{2},\dots,\lambda_{m}]^{\top}$, ${\bm{U}}=[{\bm{u}}_{1}-{\bm{u}}_{2},\dots,{\bm{u}}_{1}-{\bm{u}}_{m}]$, and ${\bm{D}}=[{\bm{g}}_{1}-{\bm{g}}_{2},\dots,{\bm{g}}_{1}-{\bm{g}}_{m}]$.
Nash-MTL [27] formulates problem (11) as a bargaining game [82, 83], where each objective is cast as a player and the utility function for each player is defined as ${\bm{g}}_{i}^{\top}{\bm{d}}$. Hence, ${\bm{d}}$ is computed by solving the following problem:
@@
FairGrad [28] extends Nash-MTL [27] by computing ${\bm{\lambda}}$ through the equation ${\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}={\bm{\lambda}}^{-1/\gamma}$, where $\gamma\geq 0$ is a constant. Unlike Nash-MTL [27], FairGrad treats the problem as a constrained nonlinear least squares problem. Similarly, UPGrad [29] calculates ${\bm{\lambda}}$ by optimizing $\min_{{\bm{\lambda}}}{\bm{\lambda}}^{\top}{\bm{G}}^{\top}{\bm{G}}{\bm{\lambda}}$, a quadratic programming (QP) problem.
Aligned-MTL [30] considers problem (11) as a linear system and aims to enhance its stability by minimizing the condition number of ${\bm{G}}$. Specifically, the Gram matrix ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ is first computed. Then, the eigenvalues $\{\sigma_{1},\dots,\sigma_{R}\}$ and eigenvectors ${\bm{V}}$ of ${\bm{C}}$ are obtained. Finally, ${\bm{\lambda}}$ is computed by:
@@
where ${{\bm{\Sigma}}}^{-1}=\text{diag}\left(\sqrt{\nicefrac{{1}}{{\sigma_{1}}}},\dots,\sqrt{\nicefrac{{1}}{{\sigma_{R}}}}\right)$.
Gradient Normalization (GradNorm) [31] aims to learn ${\bm{\lambda}}$ so that the scaled gradients have similar magnitudes. It updates ${\bm{\lambda}}$ by solving the following problem via one-step gradient descent:
@@
where $c=\frac{1}{m}\sum_{i=1}^{m}\left\|\lambda_{i}{\bm{g}}_{i}\right\|$ is the average scaled gradient norm and is treated as a constant, $r_{i}=\frac{f_{i}/f^{(0)}_{i}}{\frac{1}{m}\sum_{j=1}^{m}(f_{j}/f_{j}^{(0)})}$ is the training speed of the $i$-th objective and is also considered as a constant, and $\gamma>0$ is a hyperparameter.
@@ 3.2.2 Gradient Manipulation Methods
To address the conflicting gradient problem, gradient manipulation methods correct each objective gradient ${\bm{g}}_{i}$ to $\hat{{\bm{g}}}_{i}$ and then compute the update direction as ${\bm{d}}=\sum_{i=1}^{m}\hat{{\bm{g}}}_{i}$.
Projecting Conflicting Gradients (PCGrad) [33] considers two objective gradients ${\bm{g}}_{i}$ and ${\bm{g}}_{j}$ as conflicting if ${\bm{g}}_{i}^{\top}{\bm{g}}_{j}<0$. Then, it corrects each objective gradient by projecting it onto the normal plane of other objectives’ gradients. Specifically, at each iteration, $\hat{{\bm{g}}}_{i}$ is initialized with its original gradient ${\bm{g}}_{i}$. Then, for each $j\neq i$, if $\hat{{\bm{g}}}_{i}^{\top}{\bm{g}}_{j}<0$, $\hat{{\bm{g}}}_{i}$ is corrected as:
@@
This reduces gradient conflicts by removing the components of $\hat{{\bm{g}}}_{i}$ that oppose other objective gradients.
@@
where $\phi_{ij}$ is the cosine similarity between $\hat{{\bm{g}}}_{i}$ and ${\bm{g}}_{j}$. $\bar{\phi}_{ij}$ is initialized to 00 and updated by Exponential Moving Average (EMA), i.e., $\bar{\phi}_{ij}\leftarrow(1-\beta)\bar{\phi}_{ij}+\beta\phi_{ij}$, where $\beta$ is a hyperparameter. Note that GradVac simplifies to PCGrad when $\bar{\phi}_{ij}=0$.
@@ 3.2.3 Practical Speedup Strategy
Gradient balancing methods suffer from high computational and storage costs. Specifically, in each iteration, almost all gradient balancing methods require performing $m$ back-propagation processes to obtain all objective gradients w.r.t. the model parameter ${\bm{\theta}}\in\mathbb{R}^{d}$ and then store these gradients ${\bm{G}}\in\mathbb{R}^{d\times m}$. This can be computationally expensive when dealing with a large number of objectives or using a neural network with a large number of parameters. Moreover, many gradient balancing methods (such as MGDA [6], CAGrad [19], Nash-MTL [27], and FairGrad [28]) need to solve an optimization problem to obtain the objective weight ${\bm{\lambda}}$ in each iteration, which also increases the computational and memory costs. Hence, several strategies are proposed to alleviate this problem.
Sener and Koltun [6] use feature-level gradients (i.e., gradients w.r.t. the representations from the last shared layer) to replace the parameter-level gradients (i.e., gradients w.r.t. the shared parameters ${\bm{\theta}}$). Since the dimension of the representation is much smaller than that of the shared parameters, it can significantly reduce the computational and memory costs. This strategy is also adopted by some gradient balancing methods, such as IMTL-G [11] and Aligned-MTL [30].
Liu et al. [19] randomly select a subset of objectives to calculate the update direction in each iteration. Navon et al. [27] propose to update the objective weight ${\bm{\lambda}}$ every $\tau$ iterations instead of updating in each iteration. Although this strategy speeds up training, they observe that it may cause a noticeable drop in performance. Liu et al. [32] consider optimizing the logarithm of the MGDA objective [6] and propose a speedup strategy. Specifically, when solving problem (13) via one-step gradient descent, ${\bm{\lambda}}$ is updated as ${\bm{\lambda}}\leftarrow{\bm{\lambda}}-\eta\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}$. Note that
$\frac{1}{2}\nabla_{{\bm{\lambda}}}\left\|{\bm{G}}{\bm{\lambda}}\right\|^{2}={{\bm{G}}}^{\top}{\bm{G}}{\bm{\lambda}}={{\bm{G}}}^{\top}{\bm{d}}$
$\approx\frac{1}{\eta}\left[f^{(k)}_{1}-f^{(k+1)}_{1},\dots,f^{(k)}_{m}-f^{(k+1)}_{m}\right]^{\top},$ (24)
@@ 3.3 Discussions
Loss balancing and gradient balancing methods effectively mitigate conflicts among different objectives and achieve a Pareto-optimal solution, which is important for MOO scenarios such as multi-task learning. However, they are limited to finding a single Pareto-optimal solution and lack the ability to control the solution’s position on the Pareto front, resulting in inapplicability in some scenarios such as multi-criteria learning. In Sections 4 and 5, we introduce methods to achieve more diverse and controllable solutions.
@@
Figure 4: Distributions of solutions obtained by using representative methods: PTML [36] and EPO [37, 38] are preference vector-based methods, while GradHV [47] and MOO-SVGD [48] are methods not using preference vectors. The blue circles denote the solutions, the red curves denote the ground truth PF and the black lines denote the preference vectors.
@@ 4 Finding a Finite Set of Solutions
In some scenarios, a single solution may be insufficient to balance the objectives, as users may prefer to obtain multiple trade-off solutions and select one based on their specific needs. This section introduces algorithms to identify a finite set of solutions, providing a discrete approximation of the Pareto set. We first discuss methods based on preference vectors (Section 4.1), followed by approaches that do not require the use of preference vectors (Section 4.2). Solutions obtained by some representative methods on LSMOP1 [84] are shown in Figure 4, with additional details and more results provided in Appendix B.
4.1 Methods Based on Preference Vectors
Preference vector-based methods rely on a preference vector set $\{{\bm{\alpha}}^{(1)},\dots,{\bm{\alpha}}^{(n)}\}$. These preference vectors partition the problem into $n$ subproblems, where each subproblem involves finding the solution that corresponds to a specific preference vector in the set. By solving $n$ subproblems, these methods obtain a set of $n$ solutions.
Pareto Multi-Task Learning (PMTL) [36], inspired by the idea of decomposition-based MOO algorithms [85], incorporates preference vectors as constraints on the objectives. For the $i$-th subproblem with preference vector ${\bm{\alpha}}^{(i)}$, the objective vector ${\bm{f}}({\bm{\theta}})$ is constrained to be closer to ${\bm{\alpha}}^{(i)}$ than to the other preference vectors:
$\min_{{\bm{\theta}}\in{\mathcal{K}}\subset\mathbb{R}^{d}}~{}~{}{\bm{f}}({\bm{\theta}}):=[f_{1}({\bm{\theta}}),\dots,f_{m}({\bm{\theta}})]^{\top},$